Circular No. 8519

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005bu IN PGC 20840 Further to IAUC 8512, K. Shimasaki and W. Li report the LOSS dis-

covery of an apparent supernova on unfiltered KAIT images taken on Apr. 21.19 (mag 17.4) and 22.16 UT (mag 17.5). SN 2005bu is located at α = 7^{h}22^{m}18^{s}.21, δ = +22^{o}02′39′′.2 (equinox 2000.0), which is 7′′.3 east and 3′′.8 south of the nucleus of PGC 20840. A KAIT image taken on Apr. 1.18 showed nothing at this position (limiting mag 19.5).

COMET C/2005 G2 (SOHO) K. Battams reports that the following Marsden-group comet appeared

in C2 images as very small and faint (with no tail, as always with Marsdengroup comets), brightening slightly to mag 8.0 at 10.4R_{⊙} (Apr. 14.354 UT) via the C3 coronagraph. The initial position below continues the pattern from IAUC 8517:

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 G2 Apr. 13.685 1^{h}35.^{m}2 + 8^{o}55′ C3/2 HS 2005-G94

B. G. Marsden remarks on MPEC 2005-H24 on the possibility that C/1999 N5 separated into both C/2005 E4 (see IAUC 8594) and C/2005 G2. Z. Sekanina and P. W. Chodas (Jet Propulsion Laboratory) find that such a separation with zero relative velocity on T = 1999 July 11.2 would involve a relative deceleration of 8.6 in units of 10^{−5} the solar attraction. The best zero-velocity fit gives separation 0.93 day before perihelion and deceleration 14.6 units. For the suggested separation (IAUC 8494) of C/1999 J6 and C/1999 N5 at their previous T in Nov. 1993 (and again with zero velocity) Sekanina and Chodas find a minimum relative deceleration of 16.6 units. On MPEC 2005-H24 Marsden points out that, if C/1999 P6 and C/1999 P8/P9 (see IAUC 7863) were also produced at the Nov. 1993 separation, they should return (if they still exist) within a few days of 2005 Apr. 28 and May 18, respectively.

COMET C/2004 Q2 (MACHHOLZ) Total visual magnitude estimates: Mar. 9.98 UT, 5.7 (J. Carvajal,

Madrid, Spain, 6×30 binoculars); 18.14, 6.5 (N. Biver, Poigny-la-foret, France, 7×50 binoculars); Apr. 1.85, 7.0 (A. Kammerer, Malsch, Germany, 9×63 binoculars); 10.85, 7.2 (A. Diepvens, Balen, Belgium, 20×50 binoculars); 17.84, 7.6 (A. Baransky, Bucha, Ukraine, 10×50 binoculars).

2005 April 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green